Vaughn Corporation’s April 30 inventory was destroyed by fire. January 1 inventory was $138,300, and purchases for January through April totaled $492,100. Sales revenue for the same period was $698,900. Vaughn’s normal gross profit percentage is 35% on sales.

Using the gross profit method, estimate Vaughn’s April 30 inventory that was destroyed by fire.

Estimated ending inventory destroyed in fire
Answer #1

Answer

A Sales $698,900
B=A x 35% Gross Profits $244,615
C = A - B Cost of Goods Sold $454,285
D Beginning inventory $138,300
E Purchases $492,100
F = D+E Cost of Goods available for sale $630,400
G = F - C Estimated ending inventory destroyed in fire $176,115 Answer